# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

### v4

@@ 4.2 Overdispersion
The weights $w_{0}$ and $w_{1}$ correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that $\theta(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$. It follows that $w_{0}$ and $w_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds. We suspect this bound is conservative, however, and the typical market contains only one or two vendors.
The same behavior is exhibited by the subregion estimator of $\theta(B)$,
$\tilde{\theta}(B)=\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}$
which is asymptotically normal with mean $\theta(B)=\mu_{0}(B)$ and standard error
$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
where
$v_{0}=\dfrac{\theta(B)-M(B)}{M(B)}$
The estimator for the subregion total, $\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)$, is also asymptotically normal with mean $\tau(B)$ and standard error
$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+\left(\frac{v_{1}\,N_{1}(B)\left(N_{1}(A\backslash B)-\theta(B)\right)^{2}+v_{2}\,N_{1}(A\backslash B)\left(N_{1}(B)+\theta(B)\right)^{2}}{N_{1}(A)\,\theta(B)^{2}}\right)\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
where
$v_{1}=\dfrac{N_{1}(B)-M(B)}{M(B)+1}\qquad\text{and}\qquad v_{2}=\dfrac{N_{1}(A\backslash B)-M(A\backslash B)}{M(A\backslash B)+1}$
See Section 7.3.4 for details.
@@ 5. Discussion
We conclude by discussing the implications of our estimates. In particular, we assess the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data. Our estimates suggest that while both data sets are spatially representative, they miss the majority of street vendors. We provide several explanations for this discrepancy.
@@
Note the importance of the assumptions stated in Section 3.2 since other derivations of ratio estimation may imply smaller or larger standard errors, from which we might conclude the ACS misses a smaller or larger portion of the street vending population. For example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 300 vendors in the average market. We believe this is highly unlikely, suggesting the two estimates are inconsistent even under overdispersion.
@@
There may be several reasons why our estimates are significantly larger than the ACS. The fact that the Street Vendor Project distributed relief aid may have encouraged respondents who are unlikely to respond to the ACS. Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have covered individuals missed by the Census Bureau’s sampling frame.
Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations or concepts. For example, the ACS estimate reflects the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021. For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions. It is possible that the Street Vendor Project’s definition of a vendor is more inclusive. Street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy. In this case, either estimate may be relevant, depending on the intended use case.

### v5

@@ 5.2 Model Validation
It is possible, however, that vendors cluster within markets as described in Section 4.2. The weights $u_{0}$ and $u_{1}$ in Section 4.2 correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that $\Lambda_{0}(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$. It follows that $u_{0}$ and $u_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds. We suspect this bound is conservative, however, and the typical market contains only one or two vendors.
@@ 6. Discussion
We discuss the implications of our estimates. In particular, we assess the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data. Our estimates suggest that while both data sets are spatially representative, they miss the majority of street vendors. We provide several explanations for this discrepancy.
@@
Note the importance of the assumptions stated in Section 3.2 since deviations might suggest the ACS misses a smaller or larger portion of the street vending population. For example, if the average unpermitted or unlicensed vendor is 10 times more likely to respond than the average permitted or licensed vendor, then the ACS estimate may miss less than half of the street vending population. For another example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 300 vendors in the average market. We believe it is highly unlikely that our assumptions are violated to this extreme, suggesting the two estimates are inconsistent even under differential nonresponse and overdispersion.
@@
There may be several reasons why our estimates are significantly larger than the ACS. The fact that the Street Vendor Project distributed relief aid may have encouraged respondents who are unlikely to respond to the ACS. Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have covered individuals whose households were missed by the Census Bureau’s sampling frame. Indeed, a street vendor must be sufficiently visible while on the street in order to sell their items, while the residency of that same vendor may be more easily overlooked.
Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations or concepts. For example, the ACS estimate reflects the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021. For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions. It is possible that the Street Vendor Project’s definition of a vendor is more inclusive. Street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the “gig economy.” In this case, either estimate may be relevant, depending on the intended use case.
We believe this application highlights the value of modeling respondents as a point process. We believe our approach is relevant for a wide variety of surveys in which units cannot be chosen at random from a sampling frame. For example, individuals may be sampled on a beach or at various entrances of a park. Future work might explore additional point process models that describe the behavior of the population. For example, we have found that when vendors cluster within markets, standard errors become overdispersed. However, it is possible that vendors protect their turf, leading to underdispersion. It may even be that some vendors benefit from operating near each other, while others benefit from maintaining their distance.